### v1

# Inducing Group Fairness in LLM-Based Decisions

Abstract
Prompting Large Language Models (LLMs) has created new and interesting means for classifying textual data. While evaluating and remediating group fairness is a well-studied problem in classifier fairness literature, some classical approaches (e.g., regularization) do not carry over, and some new opportunities arise (e.g., prompt-based remediation). We measure fairness of LLM-based classifiers on a toxicity classification task, and empirically show that prompt-based classifiers may lead to unfair decisions. We introduce several remediation techniques and benchmark their fairness and performance trade-offs. We hope our work encourages more research on group fairness in LLM-based classifiers.
Inducing Group Fairness in LLM-Based Decisions
James Atwood     Preethi Lahoti     Ananth Balashankar     Flavien Prost     Ahmad Beirami Google DeepMind
1 Introduction
Large Language Models (LLMs) have shown impressive performance across tasks and are now being deployed across many high-stakes applications such as financial Wu et al. (2023) or medical Singhal et al. (2023) domains. In particular, zero-shot LLM-based classifiers Wei et al. (2022a); Anil et al. (2023) have been shown to achieve state-of-the-art performance on several natural language classification benchmarks and are being widely adopted for decision making. More recently, such classifiers are even used as a reward signal to align the model to AI Feedback Bai et al. (2022). Hence, it is important to investigate the fairness of such classifiers to different demographics.
While LLMs have been broadly studied in the generative case for diversity, stereotypes, gender bias, and toxicity Nadeem et al. (2021); Liang et al. (2021); Deshpande et al. (2023); Lahoti et al. (2023), their fairness in classification problems remains under-explored. In this paper, we investigate specifically if zero-shot and few-shot classifiers satisfy group fairness notions such as Equality of Opportunity (EO) Hardt et al. (2016), measured as the difference of false positive rates (FPR) between demographic groups.
Not surprisingly, we find that the prompted zero-shot LLM-based classifiers demonstrate a significant gap in FPR across multiple demographic groups, with Muslim and Jewish groups having 89% and 48% higher FPR as compared to the Christian group in the Civil Comments toxicity detection benchmark  Borkan et al. (2019). The gap is further increased when we compare few-shot prompting based LLM classifiers with Muslim and Jewish groups having 124% and 71% more FPR compared to the majority group.
Fairness mitigation for LLM-based classifiers is challenging due to lack of access to pretraining data and procedures. In addition, little to no fine tuning is usually done in zero-shot and few-shot classification tasks, which makes it hard to remediate through standard fairness remediation techniques for classifier fairness Hort et al. (2022).
To mitigate this gap, we benchmark the effectiveness of prompting as well as regularization-based remediation methods. In the prompting methods, we rely on instruction-following Kojima et al. (2022) and study the effectiveness of group agnostic and group aware prompts. We find that these prompt-based methods are unable to decrease the FPR gap - with Muslim and Jewish groups still having FPR about 40% higher than the majority Christian group. Then we study two in-processing Prost et al. (2019); Beutel et al. (2019) and post-processing Ţifrea et al. (2024) remediation methods by learning a softmax classification layer on the final embeddings of the LLMs, which achieve better fairness-performance trade-offs. Note that the in-processing baseline is only applicable to the fine-tuned model, and cannot be applied in the zero/few-shot classification with no training.
Contributions.
Our contributions are:
  • We benchmark the group fairness of LLM-based classifiers and show that they do not satisfy Equality of Opportunity along identity aspects such as religion, race, ethnicity, sex.
  • We introduce three remediation techniques: prompt-based, in-processing, and post-processing. We highlight that prompting-based remediations which rely on instruction tuning cannot achieve lower false positive rates, and that methods that operate on the embeddings of the LLMs (in/post-processing) are needed to achieve better fairness-performance trade-offs.
  • Our findings suggest that prompt-based methods are not effective for group fairness remediation. Additionally, in-processing remediation achieves better fairness-performance trade-offs than post-processing methods.
Related Work.
Tamkin et al. (2023) provided a method for evaluating how biased a language model may be by generating hypothetical prompts with group information. This was then used to make decisions by fitting a mixed effects model and quantified the degree of bias. The authors of the Flan-T5 model Chung et al. (2022) published group-level performance of a toxicity classifier fit to the Civil Comments Identity Borkan et al. (2019) dataset. Baldini et al. (2021) explored remediation methods for achieving equalized odds for different embedding-based classifier models. To our knowledge, this is the first paper that proposes and empirically evaluates methods for remediating zero-shot and few-shot classifiers drawn from LLMs with respect to equality of opportunity fairness. The related work for fairness in classical classification can be found in Appendix A.
2 Problem Setup
Throughout this paper, we use PaLM 2 S Anil et al. (2023) as the base model. We also use PaLM 2 L model in a model transfer experiment. We explore two settings: a ‘zero-shot’ setting, and a ‘fine-tuned’ setting where we improve the classifier performance by training a classifier, or head, on top of the final representation of the model.
For both the zero-shot and fine-tuned case we begin by wrapping the text snippet to be classified in ‘wrapper text’ that encourages the LLM to make a decision. As an example, consider a post with the text ‘first post!’ In this toy example, in order to encourage the LLM to make a decision, we wrap the text to obtain ‘is “first post!” toxic? please respond with Yes or No.’ We then send this to an LLM and ask for either a text answer (zero-shot) or model activations (fine-tuning) from the LLM.
Zero-Shot Classifier. We use a scoring-based zero-shot method for adapting LLMs to classification tasks. A schematic diagram is given in Fig. 2(a) in the appendix. Our goal is to take a text post and a classification target as input and return a distribution over negative and positive classes as output. To accomplish this, we send wrapped text to an LLM and ask it to score the ‘Yes’ and ‘No’ tokens to obtain positive and negative class scores. We then apply a softmax to obtain a probability distribution and use that distribution when making decisions. This is similar to the methodology that is followed by Chung et al. (2022).
Fine-Tuned Classifier. In the fine-tuned case, we use the activations of the final layer as input, and train a logistic regression model to predict the ground truth labels. The logistic regression model is then used for decision making. This approach can be thought of as a form of parameter efficient supervised fine tuning. A schematic diagram is given in Fig. 2(b) in the appendix.
3 Experiment Setup
Dataset and Tasks.
We use the Civil Comments Identity Borkan et al. (2019) dataset as it provides the information necessary to evaluate and remediate both group fairness and performance. For processing details, please see Appendix D.
Metrics.
We quantify the equality of opportunity fairness for each group as compared to the majority group. One metric for achieving this is the False Positive Rate ratio (FPR ratio) between the group and the corresponding majority:
$\textit{FPR}\>\textit{Ratio}=\frac{\textit{FPR}_{\textit{group}}}{\textit{FPR}_{\textit{majority}}}.$ (1)
Ideally, we want the FPR ratio to be close to one. To measure performance of the classifiers, we report Area Under the Receiver Operating Characteristic Curve (ROC AUC) over the test split which includes all demographic groups.
Group fairness remediation baselines.
We now evaluate several remediation methods. We propose a family of prompt-based methods that ask the model to be fair at different levels of specificity (in general, to demographic groups, and to a specific group). We also include in-processing and post-processing methods that are adapted from successful classical approaches:
  • Post-processing. We apply FRAPPE with an MMD regularizer base (Ţifrea et al., 2024).
  • In-processing. We apply the in-processing MMD regularization of Prost et al. (2019) to model finetuning. This baseline is inapplicable to zero/few-shot prompt-based classifiers.
  • Prompt PBF. ‘Please be as fair as possible when making a decision’ is appended to the prompt.
  • Prompt PBF2SG. ‘Please be as fair as possible when making a decision about comments about religious groups or that mention religion’ is appended to the prompt.
  • Prompt PBF2TG. ‘Please be as fair as possible when making a decision about comments that mention Judaism or Jewish people’ is appended to the prompt.
For more details about the remediation methods, please see Appendix E.
4 Experiment Results
Benchmarking group fairness without remediation.
First, we evaluate PaLM 2 S model Anil et al. (2023) with respect to equality of opportunity. These experiments follow the classification methodology described in Sec. 2 and use the Civil Comments Identity Borkan et al. (2019) dataset.
Group Zero-shot FPR Ratio Fine-tuned FPR Ratio
Muslim 1.89 2.24
Jewish 1.48 1.71
Table 1: False positive rate ratios that quantify the magnitude of violation of equality of opportunity for a classifier built on PaLM 2 S. We only include the two groups with the highest gaps. Here we used ‘Christian’ as the majority group.
The results of this evaluation for the two groups with the highest ratio gaps are given in Tab. 1 and the full table is given in Tab. 2 (in the appendix). These elevated ratios imply that there is headroom for group fairness remediation. In the rest of the experiments, we will describe remediation approaches and present their empirical performance.
Comparing the Pareto frontiers of group fairness remediation.
(a) Zero-Shot Pareto
(b) Fine-Tuned Pareto
(c) Model Transfer Pareto
Figure 1: A comparison of the Pareto frontiers of in-processing and post-processing techniques. The left plot shows the performance and fairness of zero-shot remediation and the middle plot shows fine-tuned remediation; top plots give results for the Jewish group and bottom plots gives results for the Muslim group. Prompt-based methods are single points indicated by large symbols. Note that the in-processing baseline is inapplicable in the zero-shot case. Each point for in-processing and post-processing is generated by setting different values for $\lambda$ in Equations (2) and (4) in the appendix. The bold point gives the unremediated ($\lambda=0$) case for in-processing and post-processing. The dashed and solid lines give the Pareto frontier where performance can only be gained by sacrificing fairness, for post-processing and in-processing, respectively. The right plot gives the effect of model transfer. We fit a post-processing remediation model to the PaLM 2 S model then compare the effects of applying it to PaLM 2 S (native) versus the larger PaLM 2 L model (transfer). The lines give the Pareto frontier (solid for native and dashed for transfer).
Next, we benchmark the different remediation techniques for performance vs group fairness trade-offs. The results of this are shown in Fig. 1(a). For in-processing and post-processing methods, each point in the plot is generated by varying the regularizer strength (see Appendices E.2 and E.3, respectively for more details). We observe that both methods improve fairness without significantly degrading the performance of the classifier.
We also use the activations of the final layer of the LLM to fine-tune a classifier as described in Sec. 2. The in-processed technique remediates during fine-tuning; the post-processed technique performs a separate remediation step after fine-tuning is complete. Note that this fine-tuned model generally performs better than drawing predictions directly from the LLM as described in Sec. 2. However, it requires access to a supervised dataset for fine-tuning. We compare the Pareto frontiers of fairness and performance for each remediation technique in Fig. 1(b). As before, each point in the plot is generated by varying the regularizer strength, the term that trades off between performance and fairness terms in each technique’s loss function.
There are a few takeaways from these experiments. First, the ability to fine-tune on LLM-derived activations improves the performance of the classifier as compared to zero-shot case. Second, as before, we are able to improve performance without significantly degrading the performance of the fine-tuned classifier. Third, the in-processing technique generally performs better than the post-processing technique in this environment. Finally, prompt-based methods can offer some fairness and performance benefit but are generally less general and less effective than in-processing and post-processing methods.
Transfer of remediation to an unseen model.
This experiment has a different setup than those described earlier in this section. Here, we no longer make the assumption that the LLM has ‘introspective’ abilities and thus can not return its activations to use as an embedding. Instead, we use a ‘third party’ model to embed the prefix of our query and use these embeddings for any task where they are required. More specifically, we use the Google News 128-dimensional embedding model Bengio et al. (2000). [^1] https://www.kaggle.com/models/google/nnlm.
This enables us to operate in environments where drawing embeddings from the LLM is not possible. One interesting case is where we train a post-processing remediation model on one LLM then apply to another; can we reuse the existing fairness model to improve fairness with the new model? Fig. 1(c) gives the results this model transfer learning scenario. Importantly, we find that post-processing model is still able to improve fairness when transferred (although this comes at the cost of higher performance degradation than when applied to the model that it was trained on). This suggests that we can save on inference costs by training our fairness model on a LLM with fewer parameters then apply it to a larger LLM at inference time.
Note also that the Pareto frontier achieved by the smaller PaLM 2 S model in Fig. 1(c), where ‘third party’ embeddings are used, is just slightly degraded from the Pareto frontier given in the upper plot of Fig. 1(a) where model activiations are used as embeddings. This suggests that using ‘third party’ embeddings is a reasonable choice for remediation if model activations are not available.
5 Concluding Remarks
We study the fairness of LLM-based classifiers. We identify that LLM-based classifiers may exhibit group unfairness. We introduce three remediation techniques to improve fairness while maintaining acceptable performance for LLM-based classifiers. We find that prompt-based techniques offer limited benefit and are outperformed by in-processing and post-processing techniques. We also find that the in-processing technique consistently provides favorable performance vs fairness trade-off in the fine-tuned settings.
Based on our results, if remediating a fine-tuned model, in-processing seems to be a more robust approach because it consistently provides favorable fairness versus performance tradeoffs. In other LLM-based classification settings where in-processing cannot be applied (such as zero-shot remediation and transfer tasks) the post-processing technique seems to a promising approach.
We generally find that the prompt-based remediation methods have little to no impact of prompts on fairness, while counter-intuitively, we observe that fairness-oriented prompts may slightly improve performance in some cases for the less specific ‘Please be Fair’ (PBF) and ‘Please be Fair to Super Group’ (PBF2SG) methods. This is not very surprising given that fairness is a distributional issue, and hence simple prompting may not necessarily provide the distribution matching effects that we expect from remediation.
Limitations
We would like to mention a few limitations to our work that could also be seen as opportunities for future work:
  • We find that prompt-based methods are less flexible and effective than in-processing and post-processing methods. However, we do not make an exhaustive search of possible prompts and other researchers may find prompt-based methods that work. Furthermore, it has been observed that LLM capabilities improve as LLM size increases Wei et al. (2022b), and this could have a beneficial effect on prompt-based method effectiveness as LLMs become larger and more capable.
  • Our experiments are focused on equality of opportunity (group) fairness. There is no guarantee that they will generalize to other notions of fairness, and, importantly, their application does not imply that a classifier is abstractly ‘fair.’
  • LLM-based classifier inference is very expensive when compared to simpler models, and the performance of LLM-based classifiers does not yet justify that cost (for example, our LLM-based classifiers are less effective than baseline methods given by the authors of the Civil Comments dataset paper Borkan et al. (2019)). An implicit assumption of this work is that the performance of LLM-based classifiers will improve enough over time to justify their high inference costs and become deployed systems where fairness considerations are in play.
  • We consider only one LLM (PaLM 2) and one dataset (Civil Comments Identity) in English. So it remains to be seen how much our findings generalize. Having said that, given that we already find performance disparities across subgroups in this case, the need for developing fairness remediation techniques is justifiably real.
  • We only experiment with a few-handcrafted prompts for classification, and did not compare against chain-of-thought Wei et al. (2022c), self-consistency Wang et al. (2022), and automated prompt generation Gao et al. (2021) techniques - as adapting them to induce group fairness was not trivial and is left for future work.
  • We do not compare against low-rank adaptation Hu et al. (2021), prompt-tuning Lester et al. (2021), and other parameter-efficient fine-tuning techniques Liu et al. (2022) for the in-processing method.
Appendix A Related Work in Classical Classification Fairness
There is rich literature on fairness in ‘classical’ (i.e. non-LLM-based) classification. In this paper we focus on the ‘equality of opportunity’ notion of group fairness Hardt et al. (2016) which is achieved for a group and classifier when the false positive rate (or false negative rate) of the classifier is the same for instances drawn from that group when compared with instances drawn from the majority group.
Methods for improving group fairness can generally be categorized in three main classes: pre-processing, in-processing, andpost-processing methods. Pre-processing algorithms (Feldman et al., 2015; Zemel et al., 2013; Calmon et al., 2017) transform the biased data features to a new space in which the labels and sensitive attributes are statistically independent. In-processing methods (Kamiran et al., 2010; Ristanoski et al., 2013; Quadrianto and Sharmanska, 2017; Zafar et al., 2017; Berk et al., 2017; Donini et al., 2018; Raff et al., 2018; Aghaei et al., 2019; Prost et al., 2019; Beutel et al., 2019; Grari et al., 2019; Taskesen et al., 2020; Grari et al., 2020; Cho et al., 2020; Chzhen and Schreuder, 2020; Jiang et al., 2020; Lowy et al., 2022) add a regularizer or constraint to the learning objective. Post-processing approaches (Hardt et al., 2016; Pleiss et al., 2017; Alghamdi et al., 2022; Xian et al., 2023; Ţifrea et al., 2024) improve group fairness properties by altering the final decision of the classifier. See the survey paper by Hort et al. (2022) for a more comprehensive literature survey.
Among all these classical approaches, we believe post-processing approaches are the most compatible for the LLM-based classifiers. Having said that, most post-processing approaches require access to demographic labels at test time, which is infeasible, especially for LLM-based classifiers. That is why we only focused on FRAPPÉ (Ţifrea et al., 2024) which works without such access.
Appendix B Risks
There are three risks we would like to call out:
  • Group fairness remediation improves group fairness on a training set. This may fail to generalize to a held-out set under some circumstances (for instance, if there is distributional shift).
  • Group fairness remediation improves only group fairness. Importantly, it does not guarantee improvement in other notions of fairness or make a classifier abstractly ‘fair.’
  • Group fairness remediation methods could be reversed by a malicious actor to worsen the group fairness of a classifier.
Appendix C LLM-Based Classification Methodology
(a) Zero-Shot
(b) Fine Tuning through Introspection
Figure 2: Classification flow diagrams for zero-shot and fine-tuned LLM-based classification. In both cases, decisions are encouraged via ‘text wrappers’ that nudge the LLM to make a classification decision. In the zero-shot case (Fig. 2(a)) we treat the wrapped text as a prefix and query the LLM for two postfix tokens (such as ‘Yes’ or ‘No’) that represent positive and negative decisions. We apply a softmax to these scores to obtain a probability distribution over the classification result and use this for decision making. In the fine-tuned case (Fig. 2(b)) we assume that the LLM is ‘introspective’ and can supply its activations. We instead query the LLM for the activations of its last layer to serve as an embedding. We collect those embeddings into a design matrix then fit a logistic regression model on that matrix and corresponding labels. The logistic regression model is then used for downstream decision making.
Fig. 2 gives a schematic diagram of the classification methodology described in Sec. 2. Note that we use the following wrapper text for our toxicity classifier: “‘TEXT_TO_CLASSIFY’ Categorize the comment on the basis of toxicity. If the comment is toxic (harmful, malicious, derogatory, threat, insult, identity attack, etc.) output Yes, otherwise output No.”
Appendix D Data and Data Processing
Experiments in this paper are based on the Civil Comments Identity Borkan et al. (2019) dataset. This dataset was selected because it provides the information necessary to evaluate and remediate both group fairness and performance; that is, textual data and several moderation-based labels that classifiers can be trained on and group data that can be used for evaluation and remediation with respect to group fairness.
The dataset contains 405,130 training instances, 21,293 validation instances, and 21,577 test instances. We make use of all three splits in our work.
The Civil Comments identity label and group data are represented as the proportion of raters who believe that a given text instance is an example of various moderation labels as well as various group labels. Note that the group labels correspond to the whether the content of the text is relevant to that group. Because we require binary label and group data for both remediation and evaluation, we treat any non-zero proportion of raters as a positive instance and zero values as negative instances.
Appendix E Remediation Methodology
In this section we describe our prompt-based remediation methodology and our adaptation of two ‘classical’ remediation methods to the LLM environment. This is followed by an empirical comparison of the methods.
E.1 Prompting
We explore the performance and fairness of three prompt-based methods of increasing specificity. Using a running example of remediating with respect to the Jewish group, we have:
  • Please Be Fair (PBF): ‘Please be as fair as possible when making a decision’ is appended to the prompt.
  • Please Be Fair to Super Group (PBF2SG) ‘Please be as fair as possible when making a decision about comments about religious groups or that mention religion’ is appended to the prompt.
  • Please Be Fair to This Group (PBF2TG) ‘Please be as fair as possible when making a decision about comments that mention Judaism or Jewish people’ is appended to the prompt.
This is a particularly challenging environment for prompt-based methods because we are interested in inducing group fairness, a subpopulation-level behavior, but apply the same prompt to each instance. This is also difficult to define an in-context method because it does not make sense to present an instance as being group-fair or not.
Note also that, while the methods described in the next sections provide a hyperparameter that can tune the fairness versus performance trade-off, prompt-based methods have no such capability.
E.2 In-Processing
Min Diff Prost et al. (2019) is an approach that has been successful in remediating to achieve equality of opportunity in the ‘classical’ setting. The central insight behind the Min Diff approach is that the distance between the probability of a false positive for instances from groups and majority can be included in the loss. This encourages those distributions to be closer together, which in turns pushes the false positive rates for group and majority towards each other.
In the equations that follow, let $(x,y)$ represent a feature and a label, where $x\in\mathcal{X}$, and $y\in\{0,1\}$.
We use the following loss:
$L_{\textit{IP}}=L_{\textit{CE}}(\hat{Y},Y)+\lambda\hskip 2.168pt\textit{MMD}\left(\hat{Y}|Y=0,G=0;\hat{Y}|Y=0,G=1\right).$ (2)
Where $L_{\textit{IP}}$ is the in-processing loss, $L_{\textit{CE}}$ is the usual cross entropy loss, MMD is a Maximum Mean Discrepancy kernel that gives the distance between the distributions of probability of a false positive for group and associated majority, and $\lambda$ is a parameter that trades off between the two loss terms (and thus between performance and fairness).
In order to adapt this to the LLM case, we use the Min Diff loss during fine tuning. This approach does not work in the zero-shot case where no fine tuning is performed, but the following approach does.
E.3 Post-Processing
Recent work has demonstrated how in-processing techniques can be adapted to post-processing scenarios Ţifrea et al. (2024). We leverage this approach to fit a post-processed ‘emfairening’ model. The emfairening model’s predictions are added to the unremediated models predictions in logit space:
$P_{pp}(y|x)=\sigma(z(P_{bl}(y|x))+z(P_{ef}(y|x))),$ (3)
where $\sigma$ is the logistic function, $z$ is the logit function, $P_{bl}(y|x)$ is the baseline prediction distribution, $P_{ef}(y|x)$ is the emfairening model’s distribution, and $P_{pp}(y|x)$ gives the combined post-processed distribution. The emfairening model can be trained with the following loss:
[…]
where KL is the Kullback-Leibler divergence and MMD is the Maximum Mean Discrepancy kernel. As with the Min Diff approach, the Maximum Mean Discrepancy kernel pushes the probabilities of a false positive for group and majority (and thus false positives rates for group and majority) toward each other. The KL divergence term prevents ‘catastrophic forgetting’ in the emfairening model; that is, we encourage the emfairening model to not stray too far from the performant baseline model. This ensures that we maintain acceptable classifier performance when making emfairened predictions. As with the in-processing method, $\lambda$ trades off between the two loss terms and thus dials between fairness and performance.
Note that we are free to choose the baseline model $P_{bl}(y|x)$. As such, we can apply this approach directly to the LLM in a zero-shot environment or apply it to a fine-tuned model.
Appendix F Full Benchmark Ratio Gap Results
Group Zero-Shot FPR Ratio Fine-Tuned FPR Ratio
muslim 1.89 2.24
jewish 1.48 1.71
other religion 1.40 1.32
hindu 1.39 1.46
transgender 1.24 1.63
female 1.11 1.05
black 1.06 0.90
asian 0.95 0.36
latino 0.92 0.50
other race or ethnicity 0.91 0.44
homosexual gay or lesbian 0.90 1.13
other sexual orientation 0.86 0.64
buddhist 0.75 1.08
bisexual 0.72 0.77
other gender 0.57 0.85
Table 2: False positive rate ratios that quantify the magnitude of violation of equality of opportunity for a classifier built on PaLM 2.
We only report the ratio gaps for the two groups with the highest gaps (Jewish and Muslim) in Tab. 1 in Sec. 4. The full table is given in Table 2
Appendix G Use of Scientific Artifacts
We make use of two scientific artifacts: PaLM 2 Chung et al. (2022) and the Civil Comments Identity Borkan et al. (2019) dataset. Our use of PaLM 2 is consistent with the publication guidelines of the model creators. Our use of Civil Comments is consistent with the ‘Public Domain (CC0)’ license under which it is released.

### v2

@@
-Inducing Group Fairness in LLM-Based Decisions
+Inducing Group Fairness in Prompt-Based Language Model Decisions
 Abstract
-Prompting Large Language Models (LLMs) has created new and interesting means for classifying textual data.
-While evaluating and remediating group fairness is a well-studied problem in classifier fairness literature, some classical approaches (e.g., regularization) do not carry over, and some new opportunities arise (e.g., prompt-based remediation).
-We measure fairness of LLM-based classifiers on a toxicity classification task, and empirically show that prompt-based classifiers may lead to unfair decisions.
-We introduce several remediation techniques and benchmark their fairness and performance trade-offs.
-We hope our work encourages more research on group fairness in LLM-based classifiers.
-Inducing Group Fairness in LLM-Based Decisions
-James Atwood     Preethi Lahoti     Ananth Balashankar     Flavien Prost     Ahmad Beirami Google DeepMind
+Classifiers are used throughout industry to enforce policies, ranging from the detection of toxic content to age-appropriate content filtering.
+While these classifiers serve important functions, it is also essential that they are built in ways that minimize unfair biases for users.
+One such fairness consideration is called group fairness, which desires that different sub-population of users receive equal treatment.
+This is a well-studied problem in the context of ‘classical’ classifiers.
+However, the emergence of prompt-based language model (LM) decision making has created new opportunities to solve text-based classification tasks, and the fairness properties of these new classifiers are not yet well understood.
+Further, the ‘remediation toolkit’ is incomplete for LM-based decision makers and little is understood about how to improve decision maker group fairness while maintaining classifier performance.
+This work sets out to add more tools to that toolbox.
+We introduce adaptations of existing effective approaches from the classical classifier fairness to the prompt-based classifier space.
+We also devise simple methods that take advantage of the new structure of prompt-based decision makers and operate at the prompt level.
+We compare these approaches empirically on real data.
+Our results suggest that adaptations of approaches that are effective for classical classifiers remain effective in the LM-based classifier environment.
+However, there is room for further exploration of prompt-based remediation methods (and other remediation methods that take advantage of LM structure).
+Inducing Group Fairness in Prompt-Based Language Model Decisions
+James Atwood1  Nino Scherrer2 Preethi Lahoti1Ananth Balashankar1Flavien Prost1Ahmad Beirami1 1Google DeepMind  2Google, Paradigms of Intelligence Team
 1 Introduction
-Large Language Models (LLMs) have shown impressive performance across tasks and are now being deployed across many high-stakes applications such as financial Wu et al.
+Language models (LMs) have shown impressive performance across many tasks and are now being deployed across high-stakes applications such as financial Wu et al.
 (2023) or medical Singhal et al.
 (2023) domains.
-In particular, zero-shot LLM-based classifiers Wei et al.
+In particular, zero-shot LM-based classifiers Wei et al.
 (2022a); Anil et al.
-(2023) have been shown to achieve state-of-the-art performance on several natural language classification benchmarks and are being widely adopted for decision making.
-More recently, such classifiers are even used as a reward signal to align the model to AI Feedback Bai et al.
+(2023) have achieved state-of-the-art performance on several natural language classification benchmarks and are being widely adopted for decision making.
+More recently, such classifiers are leveraged as a reward signal to align models with AI feedback Bai et al.
 (2022).
-Hence, it is important to investigate the fairness of such classifiers to different demographics.
-While LLMs have been broadly studied in the generative case for diversity, stereotypes, gender bias, and toxicity Nadeem et al.
+Hence, it is important to ask: How fair are the classification decisions made by LMs?
+In this paper, we consider two classes of LM-based classifiers: (i) prompted (“out-of-the-box”) LM classifiers and (ii) trained classifiers on top of last-layer embeddings extracted from an LM.
+We first assess whether these LM-based classifiers satisfy a widely adopted classifier group fairness notion called equal opportunity (EO) Hardt et al.
+(2016); Prost and Beutel (2020).
+EO is measured as the difference between the false positive rates (FPR) of different demographic groups, where negative outcome is considered an advantaged class.
+For example, consider a toxicity detection classifier where being labeled as toxic leads to some content moderation policy.
+It is therefore desirable for content from all demographics to be falsely marked as toxic with an equal rate.
+We find that prompted LM classifiers demonstrate a significant gap in FPR across multiple demographic groups in the Civil Comments toxicity detection benchmark  Borkan et al.
+(2019), with Muslim and Jewish groups having 89% and 48% higher FPR as compared to the Christian group.
+The gap is further increased when we compare embedding-based classifiers with Muslim and Jewish groups having 124% and 71% more FPR compared to the majority group.
+We then benchmark the effectiveness of two types of group fairness remediation techniques: (i) prompting-based and (ii) regularization-based remediation methods.
+For prompting-based methods, we study the effectiveness of different group-agnostic and group-aware fairness encouraging natural language prompts.
+In the context of regularization-based methods, we study an post-processing remediation method (Ţifrea et al., 2024) and in-processing method (Prost et al., 2019; Beutel et al., 2019).
+We find that prompt-based remediation methods are unable to decrease the FPR gap in our experiments - with Muslim and Jewish groups still having FPR about 40% higher than the Christian group.
+Contributions.
+Our contributions are:
+  • We assess the group fairness of two classes of LM-based classifiers (i.e., prompt-based and embedding-based) and show that they do not satisfy equal opportunity (EO) along identity aspects such as religion, race, ethnicity, sex.
+  • We evaluate three different remediation techniques (i.e., prompting, in-processing, and post-processing) within the two studied classes of LM-based classifiers.
+We find that prompting-based remediations fail to achieve lower false positive rates, and that regularization-based approaches achieve better fairness-performance tradeoffs across both classifier classes.
+  • We find that in-processing remediation achieves better fairness-performance trade-offs than post-processing methods, but may not be always a feasible option due to limited access to the internals of the model.
+2 Related Work
+While LMs have been broadly studied in the generative case for trustworthiness, e.g., diversity, stereotypes, gender bias, and toxicity Nadeem et al.
 (2021); Liang et al.
@@
 (2023), their fairness in classification problems remains under-explored.
-In this paper, we investigate specifically if zero-shot and few-shot classifiers satisfy group fairness notions such as Equality of Opportunity (EO) Hardt et al.
-(2016), measured as the difference of false positive rates (FPR) between demographic groups.
-Not surprisingly, we find that the prompted zero-shot LLM-based classifiers demonstrate a significant gap in FPR across multiple demographic groups, with Muslim and Jewish groups having 89% and 48% higher FPR as compared to the Christian group in the Civil Comments toxicity detection benchmark  Borkan et al.
-(2019).
-The gap is further increased when we compare few-shot prompting based LLM classifiers with Muslim and Jewish groups having 124% and 71% more FPR compared to the majority group.
-Fairness mitigation for LLM-based classifiers is challenging due to lack of access to pretraining data and procedures.
-In addition, little to no fine tuning is usually done in zero-shot and few-shot classification tasks, which makes it hard to remediate through standard fairness remediation techniques for classifier fairness Hort et al.
-(2022).
-To mitigate this gap, we benchmark the effectiveness of prompting as well as regularization-based remediation methods.
-In the prompting methods, we rely on instruction-following Kojima et al.
-(2022) and study the effectiveness of group agnostic and group aware prompts.
-We find that these prompt-based methods are unable to decrease the FPR gap - with Muslim and Jewish groups still having FPR about 40% higher than the majority Christian group.
-Then we study two in-processing Prost et al.
-(2019); Beutel et al.
-(2019) and post-processing Ţifrea et al.
-(2024) remediation methods by learning a softmax classification layer on the final embeddings of the LLMs, which achieve better fairness-performance trade-offs.
-Note that the in-processing baseline is only applicable to the fine-tuned model, and cannot be applied in the zero/few-shot classification with no training.
-Contributions.
-Our contributions are:
-  • We benchmark the group fairness of LLM-based classifiers and show that they do not satisfy Equality of Opportunity along identity aspects such as religion, race, ethnicity, sex.
-  • We introduce three remediation techniques: prompt-based, in-processing, and post-processing.
-We highlight that prompting-based remediations which rely on instruction tuning cannot achieve lower false positive rates, and that methods that operate on the embeddings of the LLMs (in/post-processing) are needed to achieve better fairness-performance trade-offs.
-  • Our findings suggest that prompt-based methods are not effective for group fairness remediation.
-Additionally, in-processing remediation achieves better fairness-performance trade-offs than post-processing methods.
-Related Work.
 Tamkin et al.
-(2023) provided a method for evaluating how biased a language model may be by generating hypothetical prompts with group information.
-This was then used to make decisions by fitting a mixed effects model and quantified the degree of bias.
+(2023) provided a method for evaluating how biased a language model may be by generating hypothetical prompts with group information and making decisions by fitting a mixed effects model.
 The authors of the Flan-T5 model Chung et al.
@@
 (2021) explored remediation methods for achieving equalized odds for different embedding-based classifier models.
-To our knowledge, this is the first paper that proposes and empirically evaluates methods for remediating zero-shot and few-shot classifiers drawn from LLMs with respect to equality of opportunity fairness.
-The related work for fairness in classical classification can be found in Appendix A.
+To our knowledge, this is the first paper that proposes and empirically evaluates methods for remediating LM-based classifiers drawn from LLMs with respect to equal opportunity fairness.
+There is also rich literature on fairness in classical (i.e.
+non-LM-based) classification.
+In this paper we focus on the equal opportunity (EO) notion of group fairness Hardt et al.
+(2016); Prost and Beutel (2020) which is achieved for a group and classifier when the false positive rate (or false negative rate) of the classifier is the same for instances drawn from that group when compared with instances drawn from the majority group.
+Methods for improving group fairness can generally be categorized in three main classes: pre-processing, in-processing, and post-processing methods.
+Pre-processing algorithms (Feldman et al., 2015; Zemel et al., 2013; Calmon et al., 2017) transform the biased data features to a new space in which the labels and sensitive attributes are statistically independent.
+In-processing methods (Kamiran et al., 2010; Ristanoski et al., 2013; Quadrianto and Sharmanska, 2017; Zafar et al., 2017; Berk et al., 2017; Donini et al., 2018; Raff et al., 2018; Aghaei et al., 2019; Prost et al., 2019; Beutel et al., 2019; Grari et al., 2019; Taskesen et al., 2020; Grari et al., 2020; Cho et al., 2020; Chzhen and Schreuder, 2020; Jiang et al., 2020; Lowy et al., 2022) add a regularizer or constraint to the learning objective.
+Post-processing approaches (Hardt et al., 2016; Pleiss et al., 2017; Alghamdi et al., 2022; Xian et al., 2023; Xian and Zhao, 2024; Ţifrea et al., 2024) improve group fairness properties by altering the final decision of the classifier.
+See the survey paper by Hort et al.
+(2022) for a more comprehensive literature survey.
+Among all these classical approaches, we believe post-processing approaches are the most compatible for the LM-based decision making.
+Having said that, most post-processing approaches require access to demographic labels at test time, which is infeasible, especially for LM-based classifiers.
+That is why we only focused on FRAPPÉ (Ţifrea et al., 2024) which works without access to demographic labels.
+(a) Prompt-based Classifier
+(b) Embedding-based Classifier
+Figure 1: Classification flow diagrams for prompt-based and embedding-based classifiers.
+Decisions are encouraged via ‘text wrappers’ that nudge the LM to make a classification decision.
+(1(a)) For prompt-based classifiers, we treat the wrapped text as a prefix and query the LM for two postfix tokens (such as ‘Yes’ or ‘No’) that represent positive and negative decisions.
+We apply a softmax to these scores to obtain a probability distribution over the classification result and use this for decision making.
+(1(b)) For embedding-based classifiers, we assume that the LM is ‘introspective’ and can supply its activations.
+We instead query the LM for the activations of its last layer to serve as an embedding.
+We collect those embeddings into a design matrix then fit a logistic regression model on that matrix and corresponding labels.
+The logistic regression model is then used for downstream decision making.
 3 Problem Setup
-Throughout this paper, we use PaLM 2 S Anil et al.
-(2023) as the base model.
-We also use PaLM 2 L model in a model transfer experiment.
-We explore two settings: a ‘zero-shot’ setting, and a ‘fine-tuned’ setting where we improve the classifier performance by training a classifier, or head, on top of the final representation of the model.
-For both the zero-shot and fine-tuned case we begin by wrapping the text snippet to be classified in ‘wrapper text’ that encourages the LLM to make a decision.
-As an example, consider a post with the text ‘first post!’ In this toy example, in order to encourage the LLM to make a decision, we wrap the text to obtain ‘is “first post!” toxic?
-please respond with Yes or No.’ We then send this to an LLM and ask for either a text answer (zero-shot) or model activations (fine-tuning) from the LLM.
-Zero-Shot Classifier.
-We use a scoring-based zero-shot method for adapting LLMs to classification tasks.
+Equal opportunity notion of group fairness.
+Consider a binary classifier where the positive outcome is considered to be a disadvantaged class, e.g., being detected as toxic.
+The widely adopted equal opportunity (EO) notion of fairness requires that different demographics are disadvantaged with the same rate, i.e., the false positive rate (FPR) of the classifier is equal for all demographics (Hardt et al., 2016; Prost and Beutel, 2020; Lowy et al., 2022; Hort et al., 2022).
+For example, this means that the content from different demographics is desired to be falsely marked as toxic at an equal rate.
+LM-based classifiers.
+We explore two classes of classifiers: (i) prompt-based LM classifiers, and (ii) trained classifier on top of extracted last-layer embeddings of LMs.
+The former does not involve any training and leverages the instruction following ability of LMs.
+In contrast, the latter requires supervised training of a classifier, and is also commonly referred to as a trained linear head on top of the last-layer representation.
+In both cases, we assume grey-box access to the LM, i.e., access to the logits (for prompt-based classifiers) or last layer activations (for embedding-based classifiers) produced by the language model prior to sampling.
+  • Prompt-based LM classifier.
+    Fig.
+1 gives a schematic diagram of the classification methodology described in Sec.
+3.
+Note that we use the following wrapper text for our toxicity classifier: We prompt an LLM with the toxicity classification task formulated as yes or no question.
+By assessing the log-probabilities of the “Yes” and “no” tokens and applying a softmax function, we are able to extract a distribution over negative and positive classes.
+A schematic diagram of the process is given in Fig.
+1(a) and our prompt is given below.
+    {TEXT_TO_CLASSIFY} Categorize the comment on the basis of toxicity. If the comment is toxic (harmful, malicious, derogatory, threat, insult, identity attack, etc.) output Yes, otherwise output No.
+  • Embedding-based classifiers.
+We prompt an LLM in the exact same way as above, but extract the final-layer activations of an LM for a training set.
+Using the collected pairs of (activation, label), we then train a logistic regression model to predict the ground truth labels.
+During inference time, we obtain the last layer embedding of the wrapped classification input (see Figure 1(b)) and feed it into the trained classifier.
+This approach can be thought of as a form of parameter efficient supervised fine-tuning.
 A schematic diagram is given in Fig.
-2(a) in the appendix.
-Our goal is to take a text post and a classification target as input and return a distribution over negative and positive classes as output.
-To accomplish this, we send wrapped text to an LLM and ask it to score the ‘Yes’ and ‘No’ tokens to obtain positive and negative class scores.
-We then apply a softmax to obtain a probability distribution and use that distribution when making decisions.
-This is similar to the methodology that is followed by Chung et al.
-(2022).
-Fine-Tuned Classifier.
-In the fine-tuned case, we use the activations of the final layer as input, and train a logistic regression model to predict the ground truth labels.
-The logistic regression model is then used for decision making.
-This approach can be thought of as a form of parameter efficient supervised fine tuning.
-A schematic diagram is given in Fig.
-2(b) in the appendix.
+1(b) in the appendix.
 4 Experiment Setup
-Dataset and Tasks.
+Models.
+We use PaLM 2 S Anil et al. (2023) as the base model, and PaLM 2 L in a model transfer experiment.
+Dataset & Tasks.
 We use the Civil Comments Identity Borkan et al.
 (2019) dataset as it provides the information necessary to evaluate and remediate both group fairness and performance.
-For processing details, please see Appendix D.
-Metrics.
+For processing details, please see Appendix C.
+Evaluation metrics.
 We quantify the equality of opportunity fairness for each group as compared to the majority group.
@@
 To measure performance of the classifiers, we report Area Under the Receiver Operating Characteristic Curve (ROC AUC) over the test split which includes all demographic groups.
-Group fairness remediation baselines.
-We now evaluate several remediation methods.
-We propose a family of prompt-based methods that ask the model to be fair at different levels of specificity (in general, to demographic groups, and to a specific group).
-We also include in-processing and post-processing methods that are adapted from successful classical approaches:
-  • Post-processing. We apply FRAPPE with an MMD regularizer base (Ţifrea et al., 2024).
-  • In-processing. We apply the in-processing MMD regularization of Prost et al. (2019) to model finetuning. This baseline is inapplicable to zero/few-shot prompt-based classifiers.
-  • Prompt PBF. ‘Please be as fair as possible when making a decision’ is appended to the prompt.
-  • Prompt PBF2SG. ‘Please be as fair as possible when making a decision about comments about religious groups or that mention religion’ is appended to the prompt.
-  • Prompt PBF2TG. ‘Please be as fair as possible when making a decision about comments that mention Judaism or Jewish people’ is appended to the prompt.
-For more details about the remediation methods, please see Appendix E.
-4 Experiment Results
-Benchmarking group fairness without remediation.
-First, we evaluate PaLM 2 S model Anil et al.
-(2023) with respect to equality of opportunity.
-These experiments follow the classification methodology described in Sec.
-2 and use the Civil Comments Identity Borkan et al.
-(2019) dataset.
-Group Zero-shot FPR Ratio Fine-tuned FPR Ratio
-Muslim 1.89 2.24
-Jewish 1.48 1.71
-Table 1: False positive rate ratios that quantify the magnitude of violation of equality of opportunity for a classifier built on PaLM 2 S.
-We only include the two groups with the highest gaps.
-Here we used ‘Christian’ as the majority group.
-The results of this evaluation for the two groups with the highest ratio gaps are given in Tab.
-1 and the full table is given in Tab.
-2 (in the appendix).
-These elevated ratios imply that there is headroom for group fairness remediation.
-In the rest of the experiments, we will describe remediation approaches and present their empirical performance.
-Comparing the Pareto frontiers of group fairness remediation.
-(a) Zero-Shot Pareto
-(b) Fine-Tuned Pareto
-(c) Model Transfer Pareto
-Figure 1: A comparison of the Pareto frontiers of in-processing and post-processing techniques.
-The left plot shows the performance and fairness of zero-shot remediation and the middle plot shows fine-tuned remediation; top plots give results for the Jewish group and bottom plots gives results for the Muslim group.
-Prompt-based methods are single points indicated by large symbols.
-Note that the in-processing baseline is inapplicable in the zero-shot case.
-Each point for in-processing and post-processing is generated by setting different values for $\lambda$ in Equations (2) and (4) in the appendix.
-The bold point gives the unremediated ($\lambda=0$) case for in-processing and post-processing.
-The dashed and solid lines give the Pareto frontier where performance can only be gained by sacrificing fairness, for post-processing and in-processing, respectively.
-The right plot gives the effect of model transfer.
-We fit a post-processing remediation model to the PaLM 2 S model then compare the effects of applying it to PaLM 2 S (native) versus the larger PaLM 2 L model (transfer).
-The lines give the Pareto frontier (solid for native and dashed for transfer).
-Next, we benchmark the different remediation techniques for performance vs group fairness trade-offs.
-The results of this are shown in Fig.
-1(a).
-For in-processing and post-processing methods, each point in the plot is generated by varying the regularizer strength (see Appendices E.2 and E.3, respectively for more details).
-We observe that both methods improve fairness without significantly degrading the performance of the classifier.
-We also use the activations of the final layer of the LLM to fine-tune a classifier as described in Sec.
-2.
-The in-processed technique remediates during fine-tuning; the post-processed technique performs a separate remediation step after fine-tuning is complete.
-Note that this fine-tuned model generally performs better than drawing predictions directly from the LLM as described in Sec.
-2.
-However, it requires access to a supervised dataset for fine-tuning.
-We compare the Pareto frontiers of fairness and performance for each remediation technique in Fig.
-1(b).
-As before, each point in the plot is generated by varying the regularizer strength, the term that trades off between performance and fairness terms in each technique’s loss function.
-There are a few takeaways from these experiments.
-First, the ability to fine-tune on LLM-derived activations improves the performance of the classifier as compared to zero-shot case.
-Second, as before, we are able to improve performance without significantly degrading the performance of the fine-tuned classifier.
-Third, the in-processing technique generally performs better than the post-processing technique in this environment.
-Finally, prompt-based methods can offer some fairness and performance benefit but are generally less general and less effective than in-processing and post-processing methods.
-Transfer of remediation to an unseen model.
-This experiment has a different setup than those described earlier in this section.
-Here, we no longer make the assumption that the LLM has ‘introspective’ abilities and thus can not return its activations to use as an embedding.
-Instead, we use a ‘third party’ model to embed the prefix of our query and use these embeddings for any task where they are required.
-More specifically, we use the Google News 128-dimensional embedding model Bengio et al.
-(2000).
-[^1] https://www.kaggle.com/models/google/nnlm.
-This enables us to operate in environments where drawing embeddings from the LLM is not possible.
-One interesting case is where we train a post-processing remediation model on one LLM then apply to another; can we reuse the existing fairness model to improve fairness with the new model?
-Fig.
-1(c) gives the results this model transfer learning scenario.
-Importantly, we find that post-processing model is still able to improve fairness when transferred (although this comes at the cost of higher performance degradation than when applied to the model that it was trained on).
-This suggests that we can save on inference costs by training our fairness model on a LLM with fewer parameters then apply it to a larger LLM at inference time.
-Note also that the Pareto frontier achieved by the smaller PaLM 2 S model in Fig.
-1(c), where ‘third party’ embeddings are used, is just slightly degraded from the Pareto frontier given in the upper plot of Fig.
-1(a) where model activiations are used as embeddings.
-This suggests that using ‘third party’ embeddings is a reasonable choice for remediation if model activations are not available.
-5 Concluding Remarks
-We study the fairness of LLM-based classifiers.
-We identify that LLM-based classifiers may exhibit group unfairness.
-We introduce three remediation techniques to improve fairness while maintaining acceptable performance for LLM-based classifiers.
-We find that prompt-based techniques offer limited benefit and are outperformed by in-processing and post-processing techniques.
-We also find that the in-processing technique consistently provides favorable performance vs fairness trade-off in the fine-tuned settings.
-Based on our results, if remediating a fine-tuned model, in-processing seems to be a more robust approach because it consistently provides favorable fairness versus performance tradeoffs.
-In other LLM-based classification settings where in-processing cannot be applied (such as zero-shot remediation and transfer tasks) the post-processing technique seems to a promising approach.
-We generally find that the prompt-based remediation methods have little to no impact of prompts on fairness, while counter-intuitively, we observe that fairness-oriented prompts may slightly improve performance in some cases for the less specific ‘Please be Fair’ (PBF) and ‘Please be Fair to Super Group’ (PBF2SG) methods.
-This is not very surprising given that fairness is a distributional issue, and hence simple prompting may not necessarily provide the distribution matching effects that we expect from remediation.
-Limitations
-We would like to mention a few limitations to our work that could also be seen as opportunities for future work:
-  • We find that prompt-based methods are less flexible and effective than in-processing and post-processing methods.
-However, we do not make an exhaustive search of possible prompts and other researchers may find prompt-based methods that work.
-Furthermore, it has been observed that LLM capabilities improve as LLM size increases Wei et al.
-(2022b), and this could have a beneficial effect on prompt-based method effectiveness as LLMs become larger and more capable.
-  • Our experiments are focused on equality of opportunity (group) fairness.
-There is no guarantee that they will generalize to other notions of fairness, and, importantly, their application does not imply that a classifier is abstractly ‘fair.’
-  • LLM-based classifier inference is very expensive when compared to simpler models, and the performance of LLM-based classifiers does not yet justify that cost (for example, our LLM-based classifiers are less effective than baseline methods given by the authors of the Civil Comments dataset paper Borkan et al.
-(2019)).
-An implicit assumption of this work is that the performance of LLM-based classifiers will improve enough over time to justify their high inference costs and become deployed systems where fairness considerations are in play.
-  • We consider only one LLM (PaLM 2) and one dataset (Civil Comments Identity) in English.
-So it remains to be seen how much our findings generalize.
-Having said that, given that we already find performance disparities across subgroups in this case, the need for developing fairness remediation techniques is justifiably real.
-  • We only experiment with a few-handcrafted prompts for classification, and did not compare against chain-of-thought Wei et al.
-(2022c), self-consistency Wang et al.
-(2022), and automated prompt generation Gao et al.
-(2021) techniques - as adapting them to induce group fairness was not trivial and is left for future work.
-  • We do not compare against low-rank adaptation Hu et al.
-(2021), prompt-tuning Lester et al.
-(2021), and other parameter-efficient fine-tuning techniques Liu et al.
-(2022) for the in-processing method.
-Appendix A Related Work in Classical Classification Fairness
-There is rich literature on fairness in ‘classical’ (i.e.
-non-LLM-based) classification.
-In this paper we focus on the ‘equality of opportunity’ notion of group fairness Hardt et al.
-(2016) which is achieved for a group and classifier when the false positive rate (or false negative rate) of the classifier is the same for instances drawn from that group when compared with instances drawn from the majority group.
-Methods for improving group fairness can generally be categorized in three main classes: pre-processing, in-processing, andpost-processing methods.
-Pre-processing algorithms (Feldman et al., 2015; Zemel et al., 2013; Calmon et al., 2017) transform the biased data features to a new space in which the labels and sensitive attributes are statistically independent.
-In-processing methods (Kamiran et al., 2010; Ristanoski et al., 2013; Quadrianto and Sharmanska, 2017; Zafar et al., 2017; Berk et al., 2017; Donini et al., 2018; Raff et al., 2018; Aghaei et al., 2019; Prost et al., 2019; Beutel et al., 2019; Grari et al., 2019; Taskesen et al., 2020; Grari et al., 2020; Cho et al., 2020; Chzhen and Schreuder, 2020; Jiang et al., 2020; Lowy et al., 2022) add a regularizer or constraint to the learning objective.
-Post-processing approaches (Hardt et al., 2016; Pleiss et al., 2017; Alghamdi et al., 2022; Xian et al., 2023; Ţifrea et al., 2024) improve group fairness properties by altering the final decision of the classifier.
-See the survey paper by Hort et al.
-(2022) for a more comprehensive literature survey.
-Among all these classical approaches, we believe post-processing approaches are the most compatible for the LLM-based classifiers.
-Having said that, most post-processing approaches require access to demographic labels at test time, which is infeasible, especially for LLM-based classifiers.
-That is why we only focused on FRAPPÉ (Ţifrea et al., 2024) which works without such access.
-Appendix B Risks
-There are three risks we would like to call out:
-  • Group fairness remediation improves group fairness on a training set. This may fail to generalize to a held-out set under some circumstances (for instance, if there is distributional shift).
-  • Group fairness remediation improves only group fairness. Importantly, it does not guarantee improvement in other notions of fairness or make a classifier abstractly ‘fair.’
-  • Group fairness remediation methods could be reversed by a malicious actor to worsen the group fairness of a classifier.
-Appendix C LLM-Based Classification Methodology
-(a) Zero-Shot
-(b) Fine Tuning through Introspection
-Figure 2: Classification flow diagrams for zero-shot and fine-tuned LLM-based classification.
-In both cases, decisions are encouraged via ‘text wrappers’ that nudge the LLM to make a classification decision.
-In the zero-shot case (Fig.
-2(a)) we treat the wrapped text as a prefix and query the LLM for two postfix tokens (such as ‘Yes’ or ‘No’) that represent positive and negative decisions.
-We apply a softmax to these scores to obtain a probability distribution over the classification result and use this for decision making.
-In the fine-tuned case (Fig.
-2(b)) we assume that the LLM is ‘introspective’ and can supply its activations.
-We instead query the LLM for the activations of its last layer to serve as an embedding.
-We collect those embeddings into a design matrix then fit a logistic regression model on that matrix and corresponding labels.
-The logistic regression model is then used for downstream decision making.
-Fig.
-2 gives a schematic diagram of the classification methodology described in Sec.
-2.
-Note that we use the following wrapper text for our toxicity classifier: “‘TEXT_TO_CLASSIFY’ Categorize the comment on the basis of toxicity.
-If the comment is toxic (harmful, malicious, derogatory, threat, insult, identity attack, etc.) output Yes, otherwise output No.”
-Appendix D Data and Data Processing
-Experiments in this paper are based on the Civil Comments Identity Borkan et al.
-(2019) dataset.
-This dataset was selected because it provides the information necessary to evaluate and remediate both group fairness and performance; that is, textual data and several moderation-based labels that classifiers can be trained on and group data that can be used for evaluation and remediation with respect to group fairness.
-The dataset contains 405,130 training instances, 21,293 validation instances, and 21,577 test instances. We make use of all three splits in our work.
-The Civil Comments identity label and group data are represented as the proportion of raters who believe that a given text instance is an example of various moderation labels as well as various group labels.
-Note that the group labels correspond to the whether the content of the text is relevant to that group.
-Because we require binary label and group data for both remediation and evaluation, we treat any non-zero proportion of raters as a positive instance and zero values as negative instances.
-Appendix E Remediation Methodology
+5 Remediation Methodology
 In this section we describe our prompt-based remediation methodology and our adaptation of two ‘classical’ remediation methods to the LLM environment.
 This is followed by an empirical comparison of the methods.
-E.1 Prompting
+5.1 Prompting
 We explore the performance and fairness of three prompt-based methods of increasing specificity. Using a running example of remediating with respect to the Jewish group, we have:
@@
 Note also that, while the methods described in the next sections provide a hyperparameter that can tune the fairness versus performance trade-off, prompt-based methods have no such capability.
-E.2 In-Processing
-Min Diff Prost et al.
-(2019) is an approach that has been successful in remediating to achieve equality of opportunity in the ‘classical’ setting.
-The central insight behind the Min Diff approach is that the distance between the probability of a false positive for instances from groups and majority can be included in the loss.
+5.2 In-Processing
+Let $(x,y)$ represent an input prompt and a label, where $x\in\mathcal{X}$, and $y\in\{0,1\}$. We use the following loss:
+$L_{\textit{IP}}=$ $L_{\textit{CE}}(\hat{Y},Y)+\lambda D\left(\hat{Y};G|Y\!=\!0\right),$ (2)
+Where $L_{\textit{IP}}$ is the in-processing loss, $L_{\textit{CE}}$ is the usual cross entropy loss for learning a classification task, and $D$ is a statistical divergence promoting the decision $\widehat{Y}$ to be independent of the sensitive attribute $G$.
+Among many choices for the statistical divergence, we focus on Min Diff Prost et al.
+(2019), which is an approach that has been successful in remediating to achieve equality of opportunity in the ‘classical’ setting in industrial settings Prost and Beutel (2020).
+The central insight behind the Min Diff approach is that the distance between the probability of a false positive for instances from groups and majority can be included as the loss.
 This encourages those distributions to be closer together, which in turns pushes the false positive rates for group and majority towards each other.
-In the equations that follow, let $(x,y)$ represent a feature and a label, where $x\in\mathcal{X}$, and $y\in\{0,1\}$.
-We use the following loss:
-$L_{\textit{IP}}=L_{\textit{CE}}(\hat{Y},Y)+\lambda\hskip 2.168pt\textit{MMD}\left(\hat{Y}|Y=0,G=0;\hat{Y}|Y=0,G=1\right).$ (2)
-Where $L_{\textit{IP}}$ is the in-processing loss, $L_{\textit{CE}}$ is the usual cross entropy loss, MMD is a Maximum Mean Discrepancy kernel that gives the distance between the distributions of probability of a false positive for group and associated majority, and $\lambda$ is a parameter that trades off between the two loss terms (and thus between performance and fairness).
-In order to adapt this to the LLM case, we use the Min Diff loss during fine tuning.
-This approach does not work in the zero-shot case where no fine tuning is performed, but the following approach does.
-E.3 Post-Processing
+In particular, Min Diff uses MMD as a Maximum Mean Discrepancy kernel that gives the distance between the distributions of probability of a false positive for group and associated majority, and $\lambda$ is a parameter that trades off between the two loss terms (and thus between performance and fairness).
+To adapt Min Diff to the LM decisions, we use the Min Diff loss during fine tuning. This approach cannot work in the zero-shot case where no fine tuning is performed.
+5.3 Post-Processing
 Recent work has demonstrated how in-processing techniques can be adapted to post-processing scenarios Ţifrea et al.
@@
 We leverage this approach to fit a post-processed ‘emfairening’ model.
-The emfairening model’s predictions are added to the unremediated models predictions in logit space:
-$P_{pp}(y|x)=\sigma(z(P_{bl}(y|x))+z(P_{ef}(y|x))),$ (3)
-where $\sigma$ is the logistic function, $z$ is the logit function, $P_{bl}(y|x)$ is the baseline prediction distribution, $P_{ef}(y|x)$ is the emfairening model’s distribution, and $P_{pp}(y|x)$ gives the combined post-processed distribution.
+The emfairening model’s predictions are added to the unremediated models predictions in logit space, i.e., for all prompts $x$ and label $y$:
+$\pi_{pp}(y|x)\propto\pi_{\text{ref}}(y|x)\pi_{\text{emf}}(y|x),$ (3)
+where $\pi_{\text{ref}}(y|x)$ is the baseline prediction distribution, $\pi_{\text{emf}}(y|x)$ is the emfairening model’s distribution, and $\pi_{pp}(y|x)$ gives the combined post-processed distribution.
 The emfairening model can be trained with the following loss:
-[…]
-where KL is the Kullback-Leibler divergence and MMD is the Maximum Mean Discrepancy kernel.
-As with the Min Diff approach, the Maximum Mean Discrepancy kernel pushes the probabilities of a false positive for group and majority (and thus false positives rates for group and majority) toward each other.
+$L_{PP}=$ $\textit{KL}\left(P_{pp}\|P_{ref}(y|x)\right)+\lambda D\left(\hat{Y};G|Y\!=\!0\right)$ (4)
+where KL is the Kullback-Leibler divergence and $D$ is a fairness promoting divergence.
+In this paper, We use the industry trusted MMD kernel Prost and Beutel (2020).
 The KL divergence term prevents ‘catastrophic forgetting’ in the emfairening model; that is, we encourage the emfairening model to not stray too far from the performant baseline model.
@@
 As with the in-processing method, $\lambda$ trades off between the two loss terms and thus dials between fairness and performance.
-Note that we are free to choose the baseline model $P_{bl}(y|x)$. As such, we can apply this approach directly to the LLM in a zero-shot environment or apply it to a fine-tuned model.
-Appendix F Full Benchmark Ratio Gap Results
-Group Zero-Shot FPR Ratio Fine-Tuned FPR Ratio
+We note that the post-processing formulation here resembles controlled decoding methods used to steer the generation of language models towards high reward outcomes Mudgal et al.
+(2024) with a reward that captures the fairness of the outcome.
+We also note that we are free to choose the baseline model $\pi_{\text{ref}}(y|x)$. As such, we can apply this approach directly to any LM in a zero-shot manner or apply it to a fine-tuned model.
+6 Experiment Results
+Group fairness without remediation.
+First, we evaluate the two classifier approaches on PaLM 2 S Anil et al.
+(2023) with respect to equality of opportunity.
+These experiments follow the methodology described in Sec.
+3 and use the Civil Comments Identity Borkan et al.
+(2019) dataset.
+Group Prompt-based Classifier Embedding-based Classifier
+Muslim 1.89 2.24
+Jewish 1.48 1.71
+Table 1: False positive rate ratios that quantify the magnitude of EO violation for a classifier built on PaLM 2 S.
+We only include the two groups with the highest gaps.
+Here we used ‘Christian’ as the majority group.
+The FPR ratios for the two groups with the highest ratio gaps are given in Tab.
+1 and the results for all groups are given in Tab.
+2 (Appendix).
+These elevated FPR ratios imply a need for group fairness remediation.
+In the remainder of this section, we will focus on remediation approaches and analyze their empirical performance vs fairness trade-offs.
+Remediation of prompt-based classifiers
+Muslim Group     Jewish Group
+(a) Prompt-based Classifiers
+(b) Embedding-based Classifiers
+(c) Model Transfer
+Figure 2: Pareto frontiers of different remediation techniques.
+The left plot shows the performance and fairness of prompt-based classifiers, and the middle plot of embedding-based classifiers.
+The unremediated classifier setting is denoted by a ‘+’ and prompting-based remediation methods are denoted by single symbols.
+Note that the in-processing baseline is inapplicable to prompt-based classifiers.
+Each point for in-processing and post-processing is generated by setting different values for $\lambda$ in Equations (2) and (4) in the appendix.
+The dashed and solid lines give the Pareto frontier where performance can only be gained by sacrificing fairness, for post-processing and in-processing, respectively.
+The right plot gives the effect of model transfer.
+We fit a post-processing remediation model to the PaLM 2 S model then compare the effects of applying it to PaLM 2 S (native) versus the larger PaLM 2 L model (transfer).
+The lines give the Pareto frontier (solid for native and dashed for transfer).
+We start with analyzing remediation techniques in prompt-based classifiers.
+Results are shown in Fig.
+2(a).
+For the post-processing method, each point in the plot is generated by varying the regularizer strength (see Appendix 5.3 for more details).
+We observe that the post-processing method improve fairness without severely degrading the performance of the classifier.
+In contrast, prompt-based remediations either increase performance while keeping the high FPR ratio, or reduce the FPR ratio at the cost of performance.
+Remediation of embedding-based classifiers.
+We compare the Pareto frontiers of fairness and performance for each remediation technique in Fig.
+2(b).
+As before, each point of regularization-based techniques in the plot is generated by varying the regularizer strength that trades off between performance and fairness terms in the objective function.
+There are a few takeaways from these experiments.
+First, embedding-based classifiers show superior performance compared to prompt-based classifiers (see Fig.
+2(a)).
+Second, as before, we observe improved group fairness for regularization-based methods without significantly degrading the performance of the classifier.
+However, the in-processing technique generally performs better than the post-processing technique in this setting.
+Finally, prompt-based remediation methods show some fairness and performance benefit but are generally less controllable and less effective than in-processing and post-processing methods.
+Transfer of remediation to an unseen model.
+In this experiment, our goal is to apply the remediation to a grey-box model which only provides access to logits (and not the last layer activations), i.e., a prompted LM.
+As such, we use the Google News 128-dimensional embedding model Bengio et al.
+(2000) [^1] https://www.kaggle.com/models/google/nnlm.
+to embed the query and use these embeddings for any subsequent remediation.
+Our setup enables us to operate in environments where drawing embeddings from the LM is not feasible.
+One interesting case is where we train a post-processing remediation model on one LM, and then apply it to remediate another LM (with grey-box access); can we reuse the existing fairness model to improve fairness with the new model?
+Fig.
+2(c) gives the results this model transfer learning scenario.
+Note also that the Pareto frontier achieved by the smaller PaLM 2 S model in Fig.
+2(c), where Google News embeddings are used, is just slightly degraded from the Pareto frontier given in the upper plot of Fig.
+2(a) where model activiations are used as embeddings, making this a promising approach for fairness remediation.
+Importantly, we find that post-processing model is still able to improve fairness when transferred (although this comes at the cost of higher performance degradation than when applied to the model that it was trained on).
+This suggests that we may be able to train universal fairness mitigation heads that could be applied to any LM with grey-box access and provide fairness benefits for classification tasks.
+7 Concluding Remarks
+Fairness is an important consideration for classifiers in industry.
+Even though the research community has made significant progress on training fair classifiers, the recent shift towards prompt-based classifiers requires exploring new fairness solutions for prompt-based decision making.
+We study the group fairness of two classes of LM-based classifiers.
+We identify that LM-based classifiers may exhibit group unfairness.
+We introduce and evaluated three remediation techniques to improve fairness while maintaining acceptable performance for LM-based classifiers.
+We find that prompt-based techniques offer limited benefit and are in general outperformed by in-processing and post-processing techniques.
+Within the scope of the evaluated prompts, we find that embedding-based classifiers exhibit superior performance compared to “out-of-the-box” prompt-based classifiers in our setup.
+In addition, we find that the in-processing method consistently provides favorable performance vs fairness trade-off on embedding-based classifiers.
+We conclude that for remediating an embedding-based classifier, in-processing is a more robust approach.
+In other LM-based classification settings where in-processing cannot be applied (prompt-based classifiers and transfer tasks) the post-processing technique provides promising results.
+We find that the prompt-based remediation methods have little to no impact of prompts on fairness, while counter-intuitively, we observe that fairness-oriented prompts may slightly improve performance in some cases for the less specific ‘Please be Fair’ (PBF) and ‘Please be Fair to Super Group’ (PBF2SG) methods.
+This is not surprising given that fairness is a distributional issue, and hence prompting may not necessarily provide the distribution matching effects that we expect from remediation.
+Appendix A Limitations
+We would like to mention a few limitations to our work that could also be seen as opportunities for future work:
+  • We find that prompting-based remediation methods are less flexible and effective than in-processing and post-processing methods.
+However, we do not make an exhaustive search of possible prompts and other researchers may find prompting-based remediation methods that work.
+Furthermore, it has been observed that the capabilities of language models improve with the model size Wei et al.
+(2022b), and this could have a beneficial effect on prompt-based method effectiveness as LMs become larger and more capable.
+  • Apart from prompting-based remediation methods, the proposed remediation techniques require grey-box access to the logits of the model prior to sampling, and may not be applicable if the model only provides black-box access.
+  • Our experiments are focused on equal opportunity (EO) notion of group fairness.
+There is no guarantee that they will generalize to other notions of fairness, and, importantly, their application does not imply that a classifier is abstractly fair, as all different notions of group fairness have their own limitations and might even be at odds with each other.
+  • LM-based classifier inference is very expensive when compared to simpler models, and the performance of LM-based classifiers does not yet justify that cost (for example, our LM-based classifiers are less effective than baseline methods given by the authors of the Civil Comments dataset paper Borkan et al.
+(2019)).
+An implicit assumption of this work is that the performance of LM-based classifiers will improve enough over time to justify their high inference costs and become deployed systems where fairness considerations are in play.
+  • We considered only one language model (PaLM 2) and one dataset (Civil Comments Identity) in English.
+So it remains to be seen how much our findings generalize.
+Having said that, given that we already find performance disparities across subgroups in this limited case, the need for developing fairness remediation techniques for LM-based decision making systems is justifiably real.
+  • We only experiment with a few-handcrafted prompts for classification, and did not compare against chain-of-thought Wei et al.
+(2022c), self-consistency Wang et al.
+(2022), and automated prompt generation Gao et al.
+(2021) techniques as adapting them to induce group fairness was not trivial and is left for future work.
+  • We do not benchmark other popular techniques, such as low-rank adaptation Hu et al.
+(2021), prompt-tuning Lester et al.
+(2021), and other parameter-efficient fine-tuning techniques Liu et al.
+(2022) for the in-processing method.
+Appendix B Risks
+There are three risks we would like to call out:
+  • Group fairness remediation improves group fairness on a training set. This may fail to generalize to a held-out set under some circumstances (for instance, if there is distributional shift).
+  • Group fairness remediation improves only group fairness. Importantly, it does not guarantee improvement in other notions of fairness or make a classifier abstractly ‘fair.’
+  • Group fairness remediation methods could be reversed by a malicious actor to worsen the group fairness of a classifier.
+Appendix C Data and Data Processing
+Experiments in this paper are based on the Civil Comments Identity Borkan et al.
+(2019) dataset.
+This dataset was selected because it provides the information necessary to evaluate and remediate both group fairness and performance; that is, textual data and several moderation-based labels that classifiers can be trained on and group data that can be used for evaluation and remediation with respect to group fairness.
+The dataset contains 405,130 training instances, 21,293 validation instances, and 21,577 test instances.
+We make use of all three splits in our work.
+The training set was used for training, validation set for classifier threshold selection, and the test set for reported results.
+The Civil Comments identity label and group data are represented as the proportion of raters who believe that a given text instance is an example of various moderation labels as well as various group labels.
+Note that the group labels correspond to the whether the content of the text is relevant to that group.
+Because we require binary label and group data for both remediation and evaluation, we treat any non-zero proportion of raters as a positive instance and zero values as negative instances.
+Appendix D Full Benchmark Ratio Gap Results
+Group Prompt-based Classifier (FPR Ratio) Embedding-based Classifier (FPR Ratio)
 muslim 1.89 2.24
@@
 other gender 0.57 0.85
-Table 2: False positive rate ratios that quantify the magnitude of violation of equality of opportunity for a classifier built on PaLM 2.
-We only report the ratio gaps for the two groups with the highest gaps (Jewish and Muslim) in Tab. 1 in Sec. 4. The full table is given in Table 2
-Appendix G Use of Scientific Artifacts
+Table 2: False positive rate ratios that quantify the magnitude of violation of equality of opportunity for an unremediated classifier built on PaLM 2 (Anil et al., 2023)
+We only report the ratio gaps for the two groups with the highest gaps (Jewish and Muslim) in Tab. 1 in Sec. 6 for the unremediated case. The full table is given in Table 2
+Appendix E Use of Scientific Artifacts
 We make use of two scientific artifacts: PaLM 2 Chung et al.